# Adversarial thermodynamics

### v1

@@ III.1 Risk aversion revisited
An analogous expression does not hold in the context of gambling, due to the fact that the utility function in that case is different. This result is complementary to the result of  [16] who quantified the dissipated fluctuating work $W_{CE}^{\mathrm{diss},\,r}$ in a thermal system which starts in equilibrium, and is driven out of it. This is the Crooks’ fluctuation relation setting. There, they find
$\beta W_{\mathrm{CE}}^{\mathrm{diss},\,r}=D_{1+r}\!\big(P_{\mathrm{F}}(w)\,\|\,P_{\mathrm{R}}(-w)\big),$ (10)
where $P_{F}(w)$ is the probability of the forward process and $P_{F}(-w)$ is the probability of the reverse process. Owing to the difference in initial starting states, it is unclear if these two expressions can be related, but it does allow and agent to decide between two different work extraction games in terms of expected utility.
@@
It is useful to study the certainty equivalent and the expected work side by side. The expectation captures the stochastic-thermodynamic viewpoint of average entropy production, while the certainty equivalent aligns with the resource-theoretic notion of generalised free energies. Analysing their joint behaviour therefore clarifies how different levels of risk aversion interpolate between the two frameworks, and provides a rationality check on the agent’s strategy.
@@
From the explanations above, we see that in the Szilard engine picture, the strategy for which the agent places the partition proportional to probability corresponds to risk neutrality $r=0$ in the economic formalism [29]. Indeed, in expected utility theory, a risk neutral individual is one for which the certainty equivalent equals the expected value of their gamble. For this risk attitude, both correspond to the thermodynamic non-equilibrium free energy.
@@
Let us now focus on strategies corresponding to $r\geq 0$, i.e. risk aversion. Here the certainty equivalent $D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})$ eq. 41 is non-negative, which follows directly from properties of the Renyi divergence in this regime [1]. Note also that since for all $\alpha\in(-\infty,\infty)$ the Renyi divergences satisfy $D_{\alpha_{1}}(P_{X}||Q_{X}^{B})\leq D_{\alpha_{2}}(P_{X}||Q_{X}^{B})$ for $\alpha_{1}<\alpha_{2}$, the certainty equivalent decreases with increasing risk aversion, as expected. Since the Rényi divergence $D_{\alpha}$ is a non-decreasing function of its order $\alpha$ [30], and since $\alpha<1$ for $r>0$, it follows that $D_{\alpha}(P_{X}\|Q_{X}^{B})\leq D_{1}(P_{X}\|Q_{X}^{B})=D(P_{X}\|Q_{X}^{B})$. The weight $(1-\alpha)$ is positive, ensuring the second term is non-negative. However, because $D_{\alpha}(P_{X}\|Q_{X}^{B})$ is strictly less than $D(P_{X}\|Q_{X}^{B})$ for non-trivial distributions ($P_{X}\neq Q_{X}^{B}$), the weighted sum $\mathbb{E}[W]$ is also strictly less than the maximum $D(P_{X}\|Q_{X}^{B})$. This shortfall compared to the non-equilibrium free energy is the price the agent pays for certainty.
@@
A subtlety arises for partition placements corresponding to risk-seeking behaviour ($r<0$), where the sign of both the certainty equivalent and the expected value may shift. Thermodynamically, risk-seeking strategies with negative expected value mark an operating point at which the agent aims to violate the second law: while rare fluctuations might yield temporary gains, they fail on average.
@@
The rationality of the agent in such regimes is described by the joint behaviour of the certainty equivalent and the expected value in the following sense: if the certainty equivalent is more negative than the expected value of the gamble, the preferences remain consistent with expected utility theory. In this case, the agent is simply expressing a preference for the risky gamble which has a non-zero probability of a gain over a certain loss. This is consistent with risk-seeking behaviour.
@@
However, risk-seeking behaviour becomes pathological if the certainty equivalent is more negative than the gamble’s worst possible outcome. In this scenario, the guaranteed loss prescribed by the certainty equivalent is strictly worse than every possible result of the gamble itself. This constitutes a violation of first-order stochastic dominance—a fundamental axiom of rational choice—as the dominated sure loss should never be chosen over the gamble that uniformly outperforms it [27, 31].
@@
Hence, for risk-seeking behaviour, we must analyse the cases 1<r<0-1<r<0 and r<1r<-1 separately. For 1<r<0-1<r<0, the certainty equivalent remains positive and greater than the non-equilibrium free energy, reflecting the fact that the agent would only give up the gamble for a guaranteed amount greater than the non-equilibrium free energy.
@@
For r<1r<-1, where the certainty equivalent flips sign, we must verify the stochastic dominance condition. In appendix G, we show that in our generalised Szilard engine, the condition for a violation of stochastic dominance in the regime where r<1r<-1 (the regime for which the sign of the certainty equivalent flips) is given by
$D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})>\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)$ (11)
@@
However, the right hand side is equal to $D_{-\infty}(P_{X}||Q_{X}^{B})$, so the condition reduces to $D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})<D_{-\infty}(P_{X}||Q_{X}^{B})$. But since for all $\alpha\in(-\infty,\infty)$ the Renyi divergences satisfy $D_{\alpha_{1}}(P_{X}||Q_{X}^{B})\leq D_{\alpha_{2}}(P_{X}||Q_{X}^{B})$ for $\alpha_{1}<\alpha_{2}$, the first-order stochastic dominance condition is never violated.
@@ IV Thermodynamic risk aversion
In the section II, we showed that the non-equilibrium free energy from thermodynamics can be understood as the maximum amount of work that can be extracted in the presence of an adversary at the level of the ensemble average. The aim of this section is to analyse the same ensemble from a finite-size regime perspective using tools developed in [17], and to prove that work extraction in the finite-size regime corresponds to the maximisation of a utility function as described in section III.
@@
Recall first that for a single round of work extraction, the extracted work in terms of the outcome $x$ of the random variable can be expressed as
$w_{X}(x)=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T.$ (12)
Again, this formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
It follows that for $n$ rounds, the work extracted is given by
$W_{n}=\sum_{i=1}^{n}\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (13)
@@
by grouping terms with the same outcome, and denoting by $N_{x}$ the number of times each outcome occurs, this can be expressed as
$W_{n}=\sum_{x}\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{N_{x}}k_{B}T$ (14)
@@
It is useful to define the type $\lambda_{x^{n}}$ of the sequence $x^{n}$. Often called its empirical distribution, the type is defined as $\lambda_{x^{n}}(x)=N_{x}/n$. Rewriting the equation above in terms of the type, we obtain an equation analogous to eq. 1 in the finite-size regime:
$\boxed{W_{n}=n(D(\lambda_{x^{n}}||Q^{B}_{X})-D(\lambda_{x^{n}}||Q^{A}_{X}))k_{B}T}$ (15)
@@
Hence, by placing the partition at position $Q_{X}^{A}=\lambda_{x^{n}}$ (when Alice’s allocation matches the type of the observed sequence), the work extracted is given by $n(D(\lambda_{x^{n}}||Q^{B}_{X}))k_{B}T$. In the finite-size regime, the work extraction problem reduces to what is essentially a decision-theoretic problem of “guessing the type” of the empirical sequence $\lambda_{x^{n}}$ that will be realised. In concurrent work [17], we show that each possible type corresponds to a particular fluctuation, with probability decaying exponentially up to sub-exponential factors as $P_{X}(\lambda_{x^{n}})\doteq\exp(-nD(\lambda_{x^{n}}\|P_{X}))$, while the potential work reward scales as $nD(\lambda_{x^{n}}\|Q_{X}^{B})$.
@@
This creates a fundamental risk-reward trade-off: Alice must select her strategy $Q_{X}^{A}$ to balance the probability of success against the amount of work extracted when successful. In appendix C, we show that the optimal strategy that maximises work extraction for a given risk tolerance (minimum success probability $\epsilon$) takes the form of an exponential family interpolating between $P_{X}$ and $Q_{X}^{B}$:
$Q_{X}^{A*}(x)=\frac{P_{X}(x)^{\mu}Q_{X}^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\mu}Q_{X}^{B}(x^{\prime})^{1-\mu}},$ (16)
@@
where the parameter $\mu$ is determined by the risk constraint. The corresponding work bound is given in terms of a Rényi divergence:
$W_{n}\geq nD_{\mu}(P_{X}\|Q_{X}^{B})k_{B}T+\frac{\mu}{1-\mu}\ln\epsilon.$ (17)
@@
The detailed derivation using the method of types [32] is based on a treatment of Kelly gambling in the finite-size regime developed in concurrent work [17]. This approach reveals that all Rényi divergences $D_{\alpha}$ acquire an operational interpretation in terms of work extraction for different risk tolerances. Previous work [20] provided operational interpretation for min and max entropies as bounds in finite-size work extraction. Here, we find similar operational interpretation for all Renyi divergences. Crucially, we also provide the work extraction strategy given the agent’s constraints. In addition, this result is identical to the expected utility formulation of the problem after identifying
$\frac{r}{1+r}=1-\frac{1}{1+r}$ (18)
with $\mu$ in eq. 16.
@@ V Conclusions
The identification of the relevant utility function shows that incorporating risk aversion into work-extraction protocols makes both stochastic sensitivity to fluctuations and the generalized free energies of thermodynamics emerge from a single principle of decision theory. In fact, the certainty equivalents defined by CARA utilities coincide exactly with Rényi divergences, demonstrating that the two modern perspectives—the stochastic description of fluctuating work and the resource-theoretic hierarchy of free energies—are not separate constraints but two ways of expressing the same decision-theoretic principles. While the generalised free energies found in the context of second laws are given in terms of the Renyi-divergence from the initial state $P_{X}$ to the equilibrium state at inverse temperature $\beta$[8], here we found that the relevant Renyi-divergences is from the state $P_{X}$ to the out of equilibrium state $Q_{X}^{B}$. Understanding the relationships between these, is an interesting open question.
Acknowledgements
MA thanks Ian Ford and Paul Skrzypczyk for useful discussions. MA was supported by the Mexican National Council of Science and Technology (CONACYT), EPSRC, the Simons Foundation It from Qubit Network, and SNF Quantum Flagship Replacement Scheme (Grant No. 215933). PhF is supported by the project FOR 2724 of the Deutsche Forschungsgemeinschaft (DFG). TS is supported by by JST ERATO Grant Number JPMJER2302, Japan. JO was supported by the Simons Foundation, It from Qubit Network, and EPSRC.
@@ Appendix A Review of Kelly betting
The Kelly [18] betting scheme is an adversarial set-up in which the gambler, Alice, allocates fractions of her wealth to the possible outcomes of a horse race, always distributing a non-zero fraction to each outcome (in order to avoid ever being completely broke). The adversary Bob, also sets odds for each possible outcome.
Suppose that the horse race is described by the random variable $X$. Bob, sets the odds $o_{x}$ to outcome $x$, and Alice allocates the fraction $f_{x}$ of her wealth to outcome $x$. With this convention for describing odds, after a round of the game, Alice’s initial wealth $W_{i}$ is multiplied by the factor $f_{x}o_{x}$:
$W_{1}=f_{x}o_{x}W_{i}$ (19)
Assuming that Alice re-invests whatever wealth she has after the previous round, her wealth evolves recursively, and at round $N$ it obeys the equation
$W_{n}=f_{x}o_{x}W_{n-1}$ (20)
@@
$\frac{W_{n}}{W_{i}}=\prod_{x}(f_{x}o_{x})^{N_{x}}$ (21)
where $N_{x}$ is the number of times that the outcome $x$ occurred.
Since the fraction by which the wealth is multiplied after every round is a random variable, it is customary to write $Q_{X}^{B}(x)=o_{x}^{-1}$ and $Q^{A}_{X}(x)=f_{x}$. Here, the subscript is there to $X$ indicate that the ratio $\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}$ is a random variable, whilst the superscripts $A$ and $B$ indicate that the fraction is chosen by Alice (resp. Bob). Altogether, the notation helps to stress the fact that even though the bets and odds are the same at every round of gambling, the amount of money made by Alice is a random variable. The expression above now becomes
$W_{n}=\prod_{x}\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{N_{x}}W_{i}$ (22)
In the limit as $n>>1$, the ratio of Alice’s initial wealth $W_{i}$ to her final wealth $W_{F}$ satisfies:
$\frac{W_{F}}{W_{i}}=\exp(n(D(P_{X}||Q_{X}^{B})-D(P_{X}||Q_{X}^{A})))$ (23)
Here, $D(P_{X}\|Q_{X})$ is the relative entropy (or Kullback-Leibler divergence). Though not a true distance metric, it is often interpreted in information theory as measuring the discrepancy between probability distributions. Since relative entropy is non-negative, Alice’s optimal strategy is to set $Q_{X}^{A}=P_{X}$, ensuring that her wealth grows at the maximum possible rate. The expression above is the well-known result by Kelly [18].

### v2

@@ III Utility maximization: connection to expected utility formulations
An analogous expression does not hold in the context of gambling, because the utility function in that case is different. A related but distinct approach was taken in Ref. Ducuara et al. (2023), which quantified the dissipated fluctuating work $W_{CE}^{\mathrm{diss},\,r}$ for a system starting in thermal equilibrium and driven out of it via a protocol satisfying the Crooks fluctuation theorem. They found
$\beta W_{CE}^{\mathrm{diss},\,r}=D_{1+r}\!\big(P_{\mathrm{F}}(w)\,\|\,P_{\mathrm{R}}(-w)\big),$ (12)
where $P_{F}(w)$ is the probability of the forward process and $P_{F}(-w)$ is the probability of the reverse process. The two results, Eqs. (11) and (12), are derived from different physical premises and are not directly comparable. Our result (11) applies to a non-equilibrium work extraction game (the adversarial Szilard engine) that does not satisfy Crooks’ theorem, as the work values are determined by the agents’ choices independently of the referee’s distribution. In contrast, the result of Ref. Ducuara et al. (2023) is rooted in the fluctuation-theoretic framework of Crooks, which relates forward and reverse processes for a system initially at equilibrium. Despite these different settings, both works highlight how Rényi divergences naturally emerge when incorporating risk sensitivity into thermodynamic reasoning.
@@
It is useful to study the certainty equivalent and the expected work side by side. The expectation captures the stochastic-thermodynamic viewpoint of average entropy production, while the certainty equivalent aligns with the resource-theoretic notion of generalized free energies. Analyzing their joint behavior therefore clarifies how different levels of risk aversion interpolate between the two frameworks, and provides a rationality check on the agent’s strategy.
@@
From the explanations above, we see that in the adversarial Szilard engine picture, the strategy for which the agent places the partition proportional to probability corresponds to risk neutrality $r=0$ in the economic formalism Abbas (2018). Indeed, in expected utility theory, a risk neutral individual is one for which the certainty equivalent equals the expected value of their gamble. For this risk attitude, both correspond to the thermodynamic nonequilibrium free energy.
@@
Let us now focus on strategies corresponding to $r\geq 0$, i.e. risk aversion. Here the certainty equivalent $D_{\frac{1}{1+r}}(P||Q^{B})$ [Eq. (11)] is non-negative, which follows directly from properties of the Rényi divergence in this regime Sagawa (2022). Note also that since the Rényi divergences satisfy van Erven and Harremoës (2014)
$D_{\alpha_{1}}(P||Q^{B})\leq D_{\alpha_{2}}(P||Q^{B})\ \ \ \text{for any}\ -\infty\leq\alpha_{1}<\alpha_{2}\leq\infty,$ (13)
the certainty equivalent decreases with increasing risk aversion, as expected. Also, since $\frac{1}{1+r}<1$ for $r>0$, it follows that $D_{\frac{1}{1+r}}(P\|Q^{B})\leq D_{1}(P\|Q^{B})=D(P\|Q^{B})$. The weight $\frac{r}{1+r}$ on the right-hand side of Eq. (9) is positive, ensuring the second term is non-negative. However, because $D_{\frac{1}{1+r}}(P\|Q^{B})$ is strictly less than $D(P\|Q^{B})$ for nontrivial distributions ($P\neq Q^{B}$), the weighted sum $W^{(r)}$ is also strictly less than the maximum $D(P\|Q^{B})$. This shortfall compared to the nonequilibrium free energy is the price the agent pays for certainty.
@@
A subtlety arises for partition placements corresponding to risk-seeking behavior ($r<0$), where the sign of both the certainty equivalent and the expected value may shift. Thermodynamically, risk-seeking strategies with negative expected value mark an operating point at which the agent aims to violate the second law: while rare fluctuations might yield temporary gains, they fail on average.
@@
The rationality of the agent in such regimes is described by the joint behavior of the certainty equivalent and the expected value in the following sense: if the certainty equivalent is more negative than the expected value of the gamble, the preferences remain consistent with expected utility theory. In this case, the agent is simply expressing a preference for the risky gamble which has a non-zero probability of a gain over a certain loss. This is consistent with risk-seeking behavior.
@@
However, risk-seeking behavior becomes pathological if the certainty equivalent is more negative than the gamble’s worst possible outcome. In this scenario, the guaranteed loss prescribed by the certainty equivalent is strictly worse than every possible result of the gamble itself. This constitutes a violation of first-order stochastic dominance — a fundamental axiom of rational choice — as the dominated sure loss should never be chosen over the gamble that uniformly outperforms it Mas-Colell et al. (1995); Ingersoll (1987). That is, the violation of the first-order stochastic dominance occurs if the certainty equivalent is more negative than the worst-case work output:
$W_{CE}<\min_{x}w(x).$ (14)
@@
Hence, for risk-seeking behavior, we must analyze the cases 1<r<0-1<r<0 and r<1r<-1 separately. For 1<r<0-1<r<0, the certainty equivalent remains positive and greater than the nonequilibrium free energy, reflecting the fact that the agent would only give up the gamble for a guaranteed amount greater than the nonequilibrium free energy.
@@
For r<1r<-1, where the certainty equivalent flips sign, we must verify the stochastic dominance condition. In Appendix F, we show that in our adversarial Szilard engine, the condition for a violation of stochastic dominance in the regime where r<1r<-1 is given by
$D_{\frac{1}{1+r}}(P||Q^{B})>\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right).$ (15)
@@
Since the right-hand side is equal to $D_{\infty}(P||Q^{B})$, the condition reduces to $D_{\frac{1}{1+r}}(P||Q^{B})>D_{\infty}(P||Q^{B})$. But from inequality (13), this never happens, and thus the first-order stochastic dominance condition is never violated.
@@ IV Thermodynamic risk aversion
In Section II, we showed that the nonequilibrium free energy from thermodynamics can be understood as the maximum amount of work that can be extracted in the presence of an adversary at the level of the ensemble average. The aim of this section is to analyze the same ensemble from a finite-size regime perspective using tools developed in Arcos et al. (2025), and to prove that work extraction in the finite-size regime corresponds to the maximization of a utility function as described in Section III.
@@
Recall first that for a single round of work extraction, the extracted work in terms of the outcome $x$ of the random variable is expressed by Eq. (5). It follows that for $n$ rounds of i.i.d., the work extracted is given by
$W_{n}(x^{n})=k_{B}T\sum_{i=1}^{n}\ln\bigg(\frac{Q^{A}(x_{i})}{Q^{B}(x_{i})}\bigg),$ (16)
@@
where $x_{i}$ is the outcome of the $i$th round and $x^{n}=(x_{1},x_{2},\cdots,x_{n})$ describes the realized sequence. By grouping terms with the same outcome, and denoting by $N_{x^{n}}(x)$ the number of times each outcome occurs in $x^{n}$, this can be expressed as
$W_{n}(x^{n})=k_{B}Tn\sum_{x}N_{x^{n}}(x)\ln\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg).$ (17)
@@
It is useful to introduce the type $\lambda_{x^{n}}$ of the sequence $x^{n}$ Csiszár (1998). Often called its empirical distribution, the type is defined as $\lambda_{x^{n}}(x)=N_{x^{n}}(x)/n$. Rewriting the equation above in terms of the type, we obtain an equation analogous to Eq. (3) in the finite-size regime:
$\boxed{W_{n}(x^{n})=k_{B}Tn(D(\lambda_{x^{n}}||Q^{B})-D(\lambda_{x^{n}}||Q^{A}))}$ (18)
@@
Hence, by placing the partition at position $Q^{A}=\lambda_{x^{n}}$ (when Alice’s allocation matches the type of the observed sequence), the work extracted is given by $k_{B}TnD(\lambda_{x^{n}}||Q^{B})$. In the finite-size regime, the work extraction problem reduces to what is essentially a decision-theoretic problem of “guessing the type” of the empirical sequence $\lambda_{x^{n}}$ that will be realized. Each possible type corresponds to a particular fluctuation with probability decaying exponentially up to sub-exponential factors, while the potential work reward scales as $nD(\lambda_{x^{n}}\|Q^{B})$. More precisely, the probability of a type is bounded as Cover and Thomas (2006)
$P(\lambda_{x^{n}})\leq\exp(-nD(\lambda_{x^{n}}\|P)).$ (19)
@@
This creates a fundamental risk-reward trade-off: Alice must select her strategy to balance the probability of success against the amount of work extracted when successful. Concretely, Alice’s strategy is to match her choice $Q^{A}$ to a particular realization of type $\lambda_{x^{n}}$ with success probability larger than a given threshold $\epsilon>0$, that is,
$P(\lambda_{x^{n}})\geq\epsilon,$ (20)
which determines the risk tolerance. From this and Eq. (19), the risk constraint can be represented as
$D(\lambda_{x^{n}}\|P)\leq\frac{1}{n}\ln\left(\frac{1}{\epsilon}\right).$ (21)
Under this constraint, we can derive the optimal strategy that maximizes work, $W_{n}(x^{n})=k_{B}TnD(\lambda_{x^{n}}||Q^{B})$. Then, the optimization problem is phrased as: Maximize the reward $D(\lambda_{x^{n}}\|Q^{B})$ with respect to $\lambda_{x^{n}}$ ($Q^{A}$) subject to the constraint (21). Note that Alice does not know which type $\lambda_{x^{n}}$ is actually realized beforehand; she bets on $Q^{A}$ and will be successful if $\lambda_{x^{n}}=Q^{A}$ is realized.
Using the standard Lagrange multiplier method, we obtain the family of optimal strategies, taking the form of an exponential family interpolating between $P$ and $Q^{B}$:
$Q^{A*,\mu}(x)=\frac{P(x)^{\mu}Q^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P(x^{\prime})^{\mu}Q^{B}(x^{\prime})^{1-\mu}},$ (22)
@@
where the parameter $0\leq\mu\leq 1$ is determined by the risk constraint $\epsilon$. See also Section 11.7 of Ref. Cover and Thomas (2006) for a formally similar optimization problem related to hypothesis testing. The corresponding work bound is given in terms of the Rényi divergence:
$W_{n}(x^{n})\geq k_{B}T\left(nD_{\mu}(P\|Q^{B})+\frac{\mu}{1-\mu}\ln\epsilon\right).$ (23)
While this is also argued in the concurrent work Arcos et al. (2025), we show its derivation in Appendix G for self-containedness. We note that there does not necessarily exist a type $\lambda_{x^{n}}$ that exactly matches $Q^{A*,\mu}$ defined by Eq. (22) for finite $n$ (especially when $Q^{A*,\mu}(x)$ is an irrational number), and thus the above argument should involve some approximation for finite $n$.
@@
Our approach reveals that all Rényi divergences $D_{\alpha}$ acquire an operational interpretation in terms of work extraction for different risk tolerances. Previous study Dahlsten et al. (2011) provided operational interpretation for min and max entropies as bounds in finite-size work extraction. Here, we find another operational interpretation for all Rényi divergences. Crucially, we also provide the work extraction strategy given the agent’s constraints. In addition, this result is identical to the expected utility formulation (8), by identifying $\mu$ to $\frac{1}{1+r}$.
@@ V Conclusions
The identification of the relevant utility function shows that incorporating risk aversion into work-extraction protocols makes both stochastic sensitivity to fluctuations and the generalized free energies of thermodynamics emerge from a single principle of decision theory. In fact, the certainty equivalents defined by CARA utilities coincide exactly with Rényi divergences, demonstrating that the two modern perspectives — the stochastic description of fluctuating work and the resource-theoretic hierarchy of free energies — are not separate constraints but two ways of expressing the same decision-theoretic principles.
Acknowledgments
MA thanks Ian Ford and Paul Skrzypczyk for useful discussions. MA was supported by the Mexican National Council of Science and Technology (CONACYT), EPSRC, the Simons Foundation It from Qubit Network, and SNF Quantum Flagship Replacement Scheme (Grant No. 215933). PhF is supported by the project FOR 2724 of the Deutsche Forschungsgemeinschaft (DFG). TS is supported by JST ERATO Grant Number JPMJER2302, Japan. JO was supported by the Simons Foundation, It from Qubit Network, and EPSRC.
@@ Appendix A Review of Kelly betting
The Kelly Kelly (1956) betting scheme is an adversarial setup in which the gambler, Alice, allocates fractions of her wealth to the possible outcomes of a horse race, always distributing a non-zero fraction to each outcome (in order to avoid ever being completely broke). The adversary Bob also sets odds for each possible outcome.
Suppose that the horse race is described by a random variable $\{x\}$. Bob, sets the odds $o_{x}$ to outcome $x$, and Alice allocates the fraction $f_{x}$ of her wealth to outcome $x$. With this convention for describing odds, after a round of the game, Alice’s initial wealth $\mathcal{W}_{i}$ is multiplied by the factor $f_{x}o_{x}$:
$\mathcal{W}_{1}=f_{x}o_{x}\mathcal{W}_{i}.$ (24)
Assuming that Alice reinvests whatever wealth she has after the previous round, her wealth evolves recursively, and at round $N$ it obeys the equation
$\mathcal{W}_{n}=f_{x}o_{x}\mathcal{W}_{n-1}.$ (25)
@@
$\frac{\mathcal{W}_{n}}{\mathcal{W}_{i}}=\prod_{x}(f_{x}o_{x})^{N_{x^{n}}(x)},$ (26)
where $N_{x^{n}}(x)$ is the number of times that the outcome $x$ occurred in the realized sequence $x^{n}=(x_{1},x_{2},\cdots,x_{n})$.
Since the fraction by which the wealth is multiplied after every round is a random variable, it is customary to write $Q^{B}(x)=o_{x}^{-1}$ and $Q^{A}(x)=f_{x}$. The superscripts $A$ and $B$ indicate that the fraction is chosen by Alice and Bob, respectively. Even though the bets and odds are the same at every round of gambling, the amount of money made by Alice is a random variable. The expression above now becomes
$\mathcal{W}_{n}=\prod_{x}\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{N_{x^{n}}(x)}\mathcal{W}_{i}.$ (27)
In the limit as $n\gg 1$, the ratio of Alice’s initial wealth $W_{i}$ to her final wealth $\mathcal{W}_{F}$ satisfies
$\frac{\mathcal{W}_{F}}{\mathcal{W}_{i}}=\exp(n(D(P||Q^{B})-D(P||Q^{A}))).$ (28)
Since the relative entropy is non-negative, Alice’s optimal strategy is to set $Q^{A}=P$, ensuring that her wealth grows at the maximum possible rate. The expression above is the well-known result by Kelly Kelly (1956).